OpenStudy (anonymous):

I WILL GIVE MEDAL AND BECOME FAN Solve the equation by completing the square. Round to the nearest tenth. x^2+8x=10 A. 1.1, 9.1 B. 1.1, -9.1 C. -1.1, 9.1 D. -1.1, -9.1 Please give me the answer and tell me how you did it! Thanks! :)

OpenStudy (anonymous):

x^2+8x+4^2 = 10 + 4^2 x^2+8x +16 = 10+16 (x+4)^2 = 26 now try to complete it by following this instruction on this site http://www.purplemath.com/modules/sqrquad.htm

OpenStudy (anonymous):

I think I understand, is it A? That's what I got...

OpenStudy (anonymous):

why do you think it is A? did you actually solved it

OpenStudy (anonymous):

I tried... haha

OpenStudy (anonymous):